OpenStudy (anonymous):

cos^2(x)-sin^2(x)=sin(x)

OpenStudy (anonymous):

i have simplified it down to \[1-2\sin^2(x)=\sin(x)\]

OpenStudy (anonymous):

is it in degrees or radians?

OpenStudy (anonymous):

radians

OpenStudy (anonymous):

pi/6,5pi/6, 3pi/2,13pi/6

OpenStudy (anonymous):

but how do you get there....

OpenStudy (anonymous):

use desmos.com/calculator type in the problem see where the lines overlap each other

OpenStudy (anonymous):

i need to show work cant